OpenStudy (anonymous):

The table shows the results of drawing one marble from a bag of colored marbles. Each marble was replaced after the outcome was recorded. If there are a total of 50 marbles in the bag, about how many of the marbles are green? Color Number of Pulls red 6 green 4 orange 7 black 3

OpenStudy (kropot72):

Total number of pulls = 20 Proportion of pulls that are green = 4/20 Approximate number of green marbles in bag = 50 * (4/20)

OpenStudy (anonymous):

Thanks you helped me alot

OpenStudy (kropot72):

You're welcome :) What is your answer?

OpenStudy (anonymous):

what do you mean

OpenStudy (kropot72):

Did you understand my post and what result did you get for the calculation?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

oh and yes i do understand

OpenStudy (kropot72):

\[50\times \frac{4}{20}=?\]

OpenStudy (anonymous):

oh 10

OpenStudy (kropot72):

You are correct now!

OpenStudy (anonymous):

yeah thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

back to school i guess

OpenStudy (kropot72):

We all slip up at times :)

OpenStudy (anonymous):

yeah. bye